OpenStudy (anonymous):

find the measure of each angle. 33. an angle with measure 8 less than the measure of it's complement

OpenStudy (king):

49 degrees

OpenStudy (anonymous):

how

OpenStudy (anonymous):

x + (x+8) = 90 2x + 8 = 90 2x = 82 x =41

OpenStudy (king):

Suppose x is the angle and y is its complement Then x+y= 90 degrees x=y-8 y-8+y=90 2y=82 y=41 x=49

OpenStudy (anonymous):

35. an angle w/ the measure of it's suplement.

OpenStudy (king):

Post as a new question